OpenStudy (anonymous):

Explain, using complete sentences, why it is important to understand any limitations on the domain and range. I WILL FAN YOU AND GIVE YOU A MEDAL <3

OpenStudy (anonymous):

Log the coordinates of the specific points in space to which your spacecraft will travel. Please remember to include the graph of your points and the lines connecting each point along with your work. Launch Area:(1,2) Point A:(2,5) Point B: (4,8) Point C: (5,12) You must show your work on each question below. Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A. (1,2) (2,4) m = y2 - y1 = 5 - 2 = 3 = 1 x2 - x1 = 2 - 1 = 3 y - 2= 1 (x - 1) y - 2 = x - 1 x + y = 1 - 2 x + y = -1 Determine the equation of the line, in point-slope form, that will get your spacecraft from Point A to Point B. (2,5) (4,8) m = y2 - y1 = 8 - 5 = 3 = 1.5 x2 - x1 = 4 - 2 = 2 y - y1 = m(x - x1) y - 5 = 1.5(x - 2) Determine the equation of the line, in slope-intercept form, that will get your spacecraft from Point B to Point C. (4,8) (6,10) m= y2 - y1 = 10 - 8 = 2 = 1 x2 - x1 = 6 - 4 = 2 y - y1 = m(x - x1) y - 8 = 1(x - 4) y - 8 = x - 4 y = x - 4 + 8 y = x + 4 In question 2, you selected one of two points (Point A or Point B) to be included in your point-slope equation. Write the point-slope form of that equation again, using the other point’s coordinates. m= y2 - y1 = 8 - 5 = 3 = 1.5 x2 - x1 = 4 - 2 = 2 y - y1 = m(x - x1) y - 8 = 1.5(x - 4) Convert the equations you arrived at in question 2 and question 4 into slope-intercept form. Question 2: y - 5 = 1.5(x - 2) y = 1.5x - 3 + 5 y = 1.5x + 2 Question 4: y - 8 = 1.5(x - 4) y = 1.5x - 6 + 8 y = 1.5x + 2 This is what the question comes after.

OpenStudy (anonymous):

Dang.

OpenStudy (anonymous):

Is this algebra?

OpenStudy (anonymous):

@Donaldweeks yes.

OpenStudy (luigi0210):

If you ask so many questions no one will want to help

hero (hero):

It is important to understand domain and range so that you can properly graph a given equation. Also, if the equation represents a real world situation such as distance, length, or volume, then understanding domain and range will enable you to properly account for such limitations so that when you are solving problems your results make sense.

OpenStudy (anonymous):

The information included in this question is the intellectual property of Florida Virtual School. Posting questions from any FLVS course and receiving answers is considered cheating.